OpenStudy (anonymous):

Select one of the factors of x*2y*3 − 11x*2y + 6y*2 − 66. a. (x*2y + 6) b. (x*2y − 6) c. (y*2 + 11) d. (y − 3)

OpenStudy (anonymous):

(x2y + 6) A :)

OpenStudy (anonymous):

heres what i had to go thurgh

OpenStudy (anonymous):

to get the answer

OpenStudy (anonymous):

(x^2)(y^3) – 11(x^2)(y) + 6(y^2) – 66 = 0 (x^2)(y^3) - 11(x^2)(y) + 6(y^2) = 66 Factor a "y" out of the left side. y[(x^2)(y^2) - 11(x^2) + 6y)] = 66 (x^2)(y^2) - 11(x^2) + 6y = 66 / y Now put the 6y on the other side. (x^2)(y^2) - 11(x^2) = (66 / y) - 6y Factor, now, the x^2 on the left side. (x^2)[y^2 - 11] = (66 / y) - 6y I would advise, now, finding a common denominator for the right side so that you can get rid of the fraction. (66 / y) - 6y = (66 / y) - ((6y^2) / y) = (66 - 6y^2) / y So now you have: (x^2)[y^2 - 11] = (66 - 6y^2) / y Divide the y^2 - 11 part off to the right side: x^2 = (66 - 6y^2) / [y * (y^2 - 11)] Notice how you can factor out a 6 in the numerator. x^2 = [6(11 - y^2)] / [y * (y^2 - 11)] Factor a negative off of the top so that you can divide out the (y^2 - 11). x^2 = [-6(y^2 - 11)] / [y * (y^2 - 11)] x^2 = -6 / y Put everything on one side so that the equation = 0. (x^2)(y) = -6 (x^2)(y) + 6 = 0. The answer is A.